OpenStudy (anonymous):

How many solutions will the following system of linear equations have? –2x + 4y = –7 y=-1/2x+5 A. 0 B. 1 C. 2 D. infinite

OpenStudy (anonymous):

As you should be able to determine by putting both equations in y=mx+b, then graph them, the lines then will intercept at exactly 1 point on the graph. This point will be the solution. There is 1 solution. Since they are linear. If, however you have non-linear systems, you may have more than 1 solution.

OpenStudy (anonymous):

Would you like me to show you how I arrived at this answer in more detail?

OpenStudy (texaschic101):

its not necessary to graph...all you have to do is compare the slopes. Do you know how to do that Brandi ?

OpenStudy (anonymous):

No @texaschic101 Can you teach me?

OpenStudy (texaschic101):

Cosmichaotic is correct in saying put the equations in y = mx + b form, where m is your slope. Once your equations are in this form, compare the slopes (the number in the m position) and the y intercepts (the number in the b position) If the slopes and the y intercepts are the same, then the lines are coincident...they lie on top of each other, and therefore have infinite slopes. If the slopes are the same, but the y intercepts are different, then the lines are parallel and have no solutions. If the slopes and the y intercepts are different, then the lines only have 1 solution.

OpenStudy (anonymous):

Nicely explained @texaschic101

OpenStudy (texaschic101):

y = -1/2x + 5 the number in the m position (the slope) is -1/2 the number in the b position (the y intercept) is 5 -2x + 4y = -7 (put this in y = mx + b form).....add 2x to both sides 4y = 2x - 7 -- divide both sides by 4 y = 2/4x - 7/4 -- reduce y = 1/2x - 7/4 the number in the m position (the slope) is 1/2 the number in the b position (the y intercept) is -7/4 now compare the slopes and the y intercepts... slopes and y intercepts are both different....this means there is only 1 solution understand ?

OpenStudy (anonymous):

Yea i get. Thank you soooo much. So much easier now.

OpenStudy (texaschic101):

no problem...you could solve it by graphing...but I only graph when I need to because I do not really like to graph...lol